# Quadratic Gaussian Splatting for Efficient and Detailed Surface Reconstruction

### v1

@@ 5 Conclusion
Discussion. Gaussian Splatting is a rapidly evolving field, with numerous works exploring its use for surface reconstruction, all developed independently from ours, such as Rade-GS [36], PGSR [4], Trim-GS [7], and MVG-Splatting [19]. Our approach differs in focus, introducing a novel representation that provides curvature information. Our contributions are orthogonal to these approaches and, in theory, could enhance their methods as well.
Limitation. Since the bounding box of QGS primitives on the image is approximated, QGS processes more tiles than typical GS methods, particularly for larger quadric surfaces, resulting in a speed disadvantage. As a surface-based representation, QGS is still limited in capturing foggy scenes or transparent objects.
@@
In this supplementary material, we provide detailed explanations of the following: (1) Calculation of the projected bounding box of primitives. (2) Details of per-tile and per-pixel sorting. (3) Solving the ray-primitive intersection equation and numerical handling. (4) Derivation of the geodesic distance formula integration. (5) Derivation of Gaussian curvature. (6) We present more qualitative results. Additionally, we include a video (QGS.mp4) that briefly explains the main workflow of QGS and showcases comparative results.
@@
Figure 9: Comparison diagram of the bounding boxes between 2DGS and QGS.
However, QGS includes concave cases, and geodesic distance complicates analytical determination of the Gaussian primitive’s rendered portion during preprocessing. Specifically, the geodesic function Equation 10 in the main text lacks an elementary inverse, so computing the equipotential $\rho(\theta_{0})=\{l^{-1}(l_{0}):l_{0}=\sigma(\theta_{0})\}$, requires numerical or approximate solutions. Even with an analytical solution for ${\rho(\theta_{0})=l^{-1}(\sigma(\theta_{0}))}$, the Gaussian distribution in non-Euclidean space means equipotential lines may not fully enclose the primitive, requiring extra boundary calculations. To simplify, we use the quadric’s circumscribing rectangular box as the 3D bounding box. Its 8 vertices are projected onto the image plane, and their 2D AABB is used as the 2D bounding box, as shown in Fig 9, columns two and three. Specifically, we first scale the geodesic function to a quadratic polynomial function:
@@
Since the quadratic curve passes through the origin, we take its positive root, representing the horizontal distance from the primitive’s vertex at a geodesic distance of $\sigma_{0}$. We compute this for the major and minor axes, i.e. $\theta=0$ and $\theta=\pi/2$, yielding $\rho_{l}$ and $\rho_{s}$, with circumscribing box radii $(\rho_{l},\rho_{s},a\max(\rho_{l}^{2},\rho_{s}^{2}))$. We then project the eight vertices of the bounding box onto the image and take the minimum 2D bounding box of the resulting polygon. Due to approximations, gaps may appear between the bounding box and the primitive as shown in the third column of Fig 9, affecting rendering speed but not quality.
@@
2DGS [12] suggests the surface lies at the median intersection point where opacity reaches 0.5. However, for surface-based representations like 2DGS and QGS, the Gaussian distribution is concentrated on the surface, making the alpha-blending order more sensitive. Most reconstruction methods [35, 36, 19, 4], including 2DGS, sort by Gaussian centroid depth, causing geometric inconsistencies, as shown in Fig 8 of the main text.
@@
$\frac{\partial{\hat{X}}}{\partial{\bar{\alpha}_{i}}}$ $=X_{i}T_{i}-\frac{\sum_{j=i+1}^{N}{X_{j}\bar{\alpha}_{j}T_{j}}}{1-\bar{\alpha}_{i}}$
$=(X_{i}-\frac{\sum_{j=i+1}^{N}{X_{j}\bar{\alpha}_{j}T_{j}}}{T_{i+1}})T_{i}$
@@
$t_{n}=\frac{-B-\mathrm{sign}(A)\cdot\sqrt{B^{2}-4AC}}{2A},$
$t_{f}=\frac{-B+\mathrm{sign}(A)\cdot\sqrt{B^{2}-4AC}}{2A}$
@@
The projection of $(\mathbf{p}_{o}-\hat{\mathbf{o}})$ onto the normal direction $\mathbf{n}$ is:
$m=\frac{\hat{\mathbf{n}}(\mathbf{p}_{o})\cdot(\mathbf{p}_{o}-\mathbf{\hat{o}})}{||\hat{\mathbf{n}}(\mathbf{p}_{o})||}$
@@
$\mathrm{cos}\theta=\frac{\mathbf{\hat{r}}\cdot\mathbf{\hat{n}}(\mathbf{p}_{o})}{||\mathbf{\hat{r}}||\cdot||\mathbf{\hat{n}}(\mathbf{p}_{o})||}$
Noting that $||\hat{\mathbf{r}}||=1$, the length of the line from the camera center along the viewing direction to the intersection with the tangent plane is given by:
$n$ $=\frac{m}{\cos\theta}=\frac{\hat{\mathbf{n}}(\mathbf{p}_{o})\cdot(\mathbf{p}_{o}-\mathbf{\hat{o}})}{\mathbf{\hat{r}}\cdot\mathbf{\hat{n}}(\mathbf{p}_{o})}$
$=-\frac{C}{B}$
@@
$n=\frac{x_{u}\times x_{v}}{||x_{u}\times x_{v}||}$ $=\frac{(-2\lambda_{1}\hat{x}_{0},-2\lambda_{2}\hat{y}_{0},1)}{\sqrt{1+4\lambda_{1}^{2}\hat{x}_{0}^{2}+4\lambda_{2}^{2}\hat{y}_{0}^{2}}}$
$x_{uu}$ $=(0,0,2\lambda_{1})$
$x_{uv}$ $=(0,0,0)$
$x_{vv}$ $=(0,0,2\lambda_{2})$
@@
Addtional Results. In this section, we present additional qualitative results of QGS in both indoor and outdoor scenarios. Fig 12 compares QGS and 2DGS on the DTU dataset, where QGS demonstrates sharper contours. QGS also achieves accurate reconstructions in indoor and outdoor scenes, as shown in Fig 13 with results from the TNT and Mip-NeRF 360 datasets. Rendering results for all three datasets are shown in Fig 14. Finally, we qualitatively tested our method on an urban scene captured from aerial view. As shown in Fig 15, QGS captures more building details than 2DGS [12] in aerial views. Additionally, most methods using the default 3DGS configuration failed to reconstruct the Courthouse scene and the urban scene, both containing nearly 1,000 images. For these two scenes, we simply doubled the training parameters, setting the total iterations to 60,000, densifying every 800 iterations from the 4,000th to the 48,000th iteration, and resetting Gaussian opacity every 6,000 iterations. For other scenes, we used the default 3DGS configuration.
Figure 12: Qualitative geometry comparison for the DTU dataset [14].
@@
Figure 15: Qualitative geometry result for the urban scene captured from aerial view, involving over 1,000 images. In the subfigure, the left shows the results of 2DGS, while the right shows the results of QGS.

### v3

@@ 5 Conclusion
Acknowledgement
This work was supported by the National Natural Science Foundation of China (Grants U22B2055 and 62273345), the Beijing Natural Science Foundation (Grant L223003), and the Key R&D Project in Henan Province (Grant 231111210300).
@@ Acknowledgement
In this supplementary material, we provide detailed explanations of the following: (1) Comparison of quadric (QGS) and disk (2DGS [16]). (2) Calculation of the projected bounding box of primitives. (3) Details of per-tile and per-pixel sorting. (4) Solving the ray-primitive intersection equation and numerical handling. (5) Derivation of the geodesic distance formula integration. (6) Derivation of Gaussian curvature. (7) We present more qualitative results.
Comparison of Disk and Quadric
We visualize the primitives of QGS (Quadric) and 2DGS (Disk), as shown in Fig 11. It can be observed that quadrics tightly conform to the surface in a curved manner, with a more noticeable fit in regions of high curvature.
(a) Mesh
(b) Disk
(c) Quadric
Figure 11: Visualizaion results on GSO dataset [9]. (a) shows the ground truth mesh, (b) displays the disk primitives, and (c) presents the quadric primitives.
@@
Figure 12: Comparison diagram of the bounding boxes between loose and tight case.
However, QGS includes concave cases, and geodesic distance complicates analytical determination of the Gaussian primitive’s rendered portion during preprocessing. Specifically, the geodesic function Equation 10 in the main text lacks an elementary inverse, so computing the equipotential $\rho(\theta_{0})=\{l^{-1}(l_{0}):l_{0}=\sigma(\theta_{0})\}$, requires numerical or approximate solutions. Moreover, even with an analytical solution for $\rho(\theta_{0})=l^{-1}(\sigma(\theta_{0}))$}, the Gaussian distribution in non-Euclidean space means equipotential lines may not fully enclose the primitive, requiring extra boundary calculations. A straightforward approach is to use the projection of the quadric’s 3D bounding box as the 2D bounding box, as shown in Fig 12 (a). Its eight vertices are projected onto the image plane, and their 2D axis-aligned bounding box (AABB) is used as the 2D bounding box, as illustrated in Fig 12 (c). However, this loose case significantly enlarges the 2D bounding box, leading to a decrease in efficiency. To address this, we construct a tighter bounding box using the quadric’s tangent planes, as shown in Fig 12 (b). Specifically, we first scale the geodesic function to a quadratic polynomial function:
@@
Since the Equation 20 through the origin, we take its positive root, representing the horizontal distance from the primitive’s vertex at a geodesic distance of $\sigma_{0}(\theta_{0})$. We compute this for the major and minor axes, i.e. $\theta=0$ and $\theta=\pi/2$, yielding $\rho_{1}$ and $\rho_{2}$, and obtain four points: $(\pm\rho_{1},0,\rho_{1}^{2}),(0,\pm\rho_{2},\rho_{2}^{2})$. We then compute the intersection of the tangent planes at these four points with the plane $z=0$.
$x_{1}=\pm\frac{\rho_{1}s_{1}-\rho_{1}^{2}}{s_{1}},\;\;x_{2}=\pm\frac{\rho_{2}s_{2}-\rho_{2}^{2}}{s_{2}}$ (22)
At this stage, we obtain a 3D truncated pyramid with top dimensions $\rho_{1}\times\rho_{2}$, bottom dimensions $x_{1}\times x_{2}$, and height $\max(\rho_{1}^{2},\rho_{2}^{2})$. We then project its eight vertices onto the image plane and determine the minimum 2D bounding box enclosing the resulting polygon as shown in Fig 12 (d).
Mip-NeRF360 TNT
Storage FPS Training Time Storage FPS
2DGS 556MB 15.34 1h5min 218MB 31.66
GOF 825MB 7.61 2h12min 367MB 13.45
QGS 688MB 7.92 1h48min 274MB 13.27
QGS w/ TB 641MB 14.15 1h13min 263MB 25.36
Table 6: Speed and storage comparison on the Mip-NeRF 360 [2] and TNT [22] datasets. With a tighter bounding box, QGS achieves a noticeable speed improvement.
By introducing a tighter bounding box and optimizing global memory access on the GPU, QGS achieves twice the speed. We denote the accelerated QGS as QGS w/ TB and compare the storage and speed of different methods on the Mip-NeRF 360 and TNT datasets as shown in Table 6.
@@
2DGS [16] suggests the surface lies at the median intersection point where opacity reaches 0.5. However, for surface-based representations like 2DGS and QGS, the Gaussian distribution is concentrated on the surface, making the alpha-blending order more sensitive.
@@
$\frac{\partial{\hat{X}}}{\partial{\bar{\alpha}_{i}}}$ $=(X_{i}-\frac{\sum_{j=i+1}^{N}{X_{j}\bar{\alpha}_{j}T_{j}}}{T_{i+1}})T_{i}$
@@
$t_{n,f}=\frac{-B\pm\mathrm{sign}(A)\cdot\sqrt{B^{2}-4AC}}{2A}$
@@
[…]
@@
[…]
Noting that $\hat{\mathbf{r}}||=1$, the length of the line from the camera center along the viewing direction to the intersection with the tangent plane is given by:
$n$ $=\frac{m}{\cos\theta}=\frac{\hat{\mathbf{n}}(\mathbf{p}_{o})\cdot(\mathbf{p}_{o}-\mathbf{\hat{o}})}{\mathbf{\hat{r}}\cdot\mathbf{\hat{n}}(\mathbf{p}_{o})}=-\frac{C}{B}$
@@
[…]
$x_{uu}$ $=(0,0,2\lambda_{1}),\;x_{uv}=(0,0,0),\;x_{vv}=(0,0,2\lambda_{2})$
@@
Addtional Results. In this section, we present additional qualitative results of QGS in both indoor and outdoor scenarios. Figure 15 presents a comparison of QGS with 2DGS [16], GOF [50], and PGSR [5] on the DTU dataset, illustrating that QGS captures more geometric details. QGS also achieves accurate reconstructions in indoor and outdoor scenes, as shown in Fig 16 with results from the TNT and Mip-NeRF 360 datasets. Rendering results for all three datasets are shown in Fig 17. Finally, we qualitatively tested our method on an urban scene captured from aerial view. As shown in Fig 18, QGS captures more building details than 2DGS [16] in aerial views.
scan37
scan55
scan63
scan69
scan110
Reference Image
2DGS [16]
GOF [50]
PGSR [5]
Ours
Figure 15: Qualitative geometric reconstruction comparisons on the DTU dataset [19]. Our method achieves reconstructions of higher quality and greater detail.
@@
Figure 18: Qualitative geometry result for the urban scene captured from aerial view, involving over 1,000 images. In the subfigure, the left shows the results of 2DGS [16], while the right shows the results of QGS.